OpenStudy (anonymous):

@Loser66 this will take a long time so be prepared !

OpenStudy (loser66):

hahaha, why me? I don't have enough time.

OpenStudy (magbak):

Hey @Loser66

OpenStudy (anonymous):

:( i need to prepare for a big test

OpenStudy (loser66):

nice to see you again @magbak

OpenStudy (anonymous):

and you're the best person to explain things to me

OpenStudy (loser66):

@plohrr just post

OpenStudy (magbak):

Nice to see you too

OpenStudy (anonymous):

its not really a question, i just need to know everything about factoring .

OpenStudy (loser66):

What don't you understand?

OpenStudy (anonymous):

factoring polynomials, basically everything

OpenStudy (loser66):

please, specific

OpenStudy (anonymous):

polynomials

OpenStudy (loser66):

post the problem you stuck

OpenStudy (anonymous):

for example, x^2 + x + 1. that is the format. how would i begin to factor this ?

OpenStudy (loser66):

that polynomial is irreducible

OpenStudy (anonymous):

?

OpenStudy (loser66):

cannot factor

OpenStudy (anonymous):

what is a factorable polynomial ?

OpenStudy (loser66):

x^2 + 4x + 4 ok?

OpenStudy (anonymous):

yes .

OpenStudy (loser66):

4 = ? *?

OpenStudy (anonymous):

2 x 2

OpenStudy (loser66):

you have options: 1*4 = 4 2*2 = 4

OpenStudy (loser66):

that is the last 4 in x^2 + 4x + 4 ( this guy)

OpenStudy (loser66):

now you test, the first option 1*4 =4 and 1+4 = 4 or not? (this 4 is the first 4 in x^2 +4x (this guy)

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

the second option works

OpenStudy (loser66):

ye, @nikato help this guy, please. there is someone answer my post. I must be there to get help. Please

OpenStudy (loser66):

@nikato Are you willing?

OpenStudy (loser66):

ok, @plohrr second question

OpenStudy (anonymous):

can you factor functions ?

OpenStudy (loser66):

yes, if it is possible

OpenStudy (anonymous):

what is a factorable one

OpenStudy (loser66):

pick from your review sheet/book

OpenStudy (anonymous):

x^2 - 6x + 13 = 0